# Angular-momentum pairs in spherical systems: applications to the Galactic centre

### v1

@@ Appendix C VRR energy for circular orbits and equations of motion
$S^{\prime}_{ij}(x,z)=$ $\sum_{\ell=2}^{\infty}[P_{\ell}(0)]^{2}P^{\prime}_{\ell}(x)z^{\ell}\approx\left(\frac{1}{\pi}-\frac{1}{4}\right)P^{\prime}_{2}(x)z^{2}-\left(\frac{1}{2\pi}-\frac{9}{64}\right)P^{\prime}_{4}(x)z^{4}+\frac{z(1+X)}{\pi(1-xz+X)X}-\frac{z(1+Y)}{\pi(1+xz+Y)Y}.$ (72)
@@
This is equation (20), which we utilise this expression throughout the paper to estimate the torques between the initially coupled angular momentum binaries (two nearly coplanar circular rings).
@@ C.1 Asymptotic equations of motion with corrections
Here $\mathcal{J}_{ij\ell}^{\rm exact}-\mathcal{J}_{ij\ell}^{\rm asymp}\propto\ell^{-3}$ for large $\ell$ which converges quickly, implying that this sum may be truncated at low orders. Taking the leading order term only, i.e. the asymptotic sum ($\propto\cot\theta$), is generally correct to within a $40\%$ error, including the first correction ($\ell=2$, $\propto\cos\theta$) is correct to $10\%$, and with the second term ($\ell=4$, $\propto\cos^{3}\theta$) it is correct to $2\%$. However for small angles the accuracy scales with $\sin\theta$. The expression $k_{ij}$ depends on the function $I^{(2)}$ (given below in Appendix C.2) and, in general case, can be evaluated numerically, but for the case of a circular ring embedded within an elliptical one, it simplifies analytically, as we show below. Additionally, $\mathcal{J}_{ij\ell}^{\rm exact}$ depends on the coefficient $s_{ijl}$ given by Eq. (4) which can be evaluated numerically, but simplifies for some particular cases (see Appendix B of Kocsis & Tremaine 2015). And again, for the case of a circular ring within an elliptical, one can get rid of the outer integral.
@@ C.2 Endpoint expansion of $I^{(2)}$ for $c\to b^{+}$
For a circular orbit embedded within an elliptical one, we can simplify $I^{(2)}(a,b,c,d)$, since in this case $b=c$. We have
@@ Appendix D Numerical simulations
We utilise 2 different codes: N-Ring (which solves equation of motion 7, Kocsis & Tremaine 2015) to check the analytical calculations, and $\phi$-CPU to run 4-body simulations which solve Newtonian equation of motion, which we describe briefly below.
@@ D.1 N-Ring
N-Ring constructs a global symplectic map by composing all $N(N-1)/2$ pair operators $\mathcal{O}_{ij}(\Delta t)=\exp(\Delta t\,\mathcal{G}_{ij})$ in a time-reversible order, yielding a second-order scheme that can be promoted to eighth order through triple-jump factorisation. A multilevel block-time-step refinement decreases the step size only for the innermost or most strongly-coupled stars, while maintaining symplecticity, and a binary-tree tiling lets non-overlapping pairs update simultaneously, giving $\mathcal{O}(N^{2}/P)+2P$ scaling on $P$ processors. The result is an exactly orbital-energy-conserving and angular-momentum-vector conserving integrator whose dominant cost is $N^{2}$ pair rotations, but whose accuracy can be tuned by the time-step hierarchy and the chosen operator-splitting order. Details on N-Ring may be found in Kocsis & Tremaine (2015).

### v2

@@ Appendix C VRR energy for circular orbits and equations of motion
$S^{\prime}_{ij}(x,z)=$ $\sum_{\ell=2}^{\infty}[P_{\ell}(0)]^{2}P^{\prime}_{\ell}(x)z^{\ell}\approx-\left(\frac{1}{\pi}-\frac{1}{4}\right)P^{\prime}_{2}(x)z^{2}-\left(\frac{1}{2\pi}-\frac{9}{64}\right)P^{\prime}_{4}(x)z^{4}+\frac{z(1+X)}{\pi(1-xz+X)X}-\frac{z(1+Y)}{\pi(1+xz+Y)Y}.$ (85)
@@
This is equation (21), which we utilise throughout the paper to estimate the torques between the initially coupled angular momentum binaries (two nearly coplanar circular rings).
@@ C.1 Asymptotic equations of motion with corrections
Here $\mathcal{J}_{ij\ell}^{\rm exact}-\mathcal{J}_{ij\ell}^{\rm asymp}\propto\ell^{-3}$ for large $\ell$ which converges quickly, implying that this sum may be truncated at low orders. Taking the leading order term only, i.e. the asymptotic sum ($\propto\cot\theta$), is generally correct to within a $40\%$ error, including the first correction ($\ell=2$, $\propto\cos\theta$) is correct to $10\%$, and with the second term ($\ell=4$, $\propto\cos^{3}\theta$) it is correct to $2\%$. However for small angles the accuracy scales with $\sin\theta$. The expression $k_{ij}$ depends on the function $I^{(2)}$ (given below in Appendix C.2) and, in general case, can be evaluated numerically, but for the case of a circular ring embedded within an elliptical one, it simplifies analytically, as we show below. Additionally, $\mathcal{J}_{ij\ell}^{\rm exact}$ depends on the coefficient $s_{ij\ell}$ given by Eq. (4) which can be evaluated numerically, but simplifies for some particular cases (see Appendix B of Kocsis and Tremaine 2015). And again, for the case of a circular ring within an elliptical, one can get rid of the outer integral.
@@ C.2 Endpoint expansion of $I^{(2)}$ for $c\to b^{+}$
For a circular orbit embedded within an elliptical one, we can simplify $I^{(2)}(a,b,c,d)$, since in this case $b=c$. Embedded orbits (rings) are a subclass of overlapping orbits in which one ring is radially embedded within the other, such that $r_{p,1}<r_{p,2}<r_{a,2}<r_{a,1}$. If the embedded ring is circular, then $r_{p,2}=r_{a,2}$, corresponding to the limit $c\to b$. In this case, the other ring must be elliptical.
We have
@@ Appendix D Numerical simulations
We utilise 2 different codes: N-Ring (which solves equation of motion 7, Kocsis and Tremaine 2015) to check the analytical calculations, and $\phi$-CPU to run 4-body simulations which solve Newtonian equation of motion, which we describe briefly below.
@@ D.1 N-Ring
N-Ring constructs a global symplectic map by composing all N(N1)/2N(N-1)/2 pair operators $\mathcal{O}_{ij}(\Delta t)=\exp(\Delta t\,\mathcal{G}_{ij})$ in a time-reversible order, yielding a second-order scheme that can be promoted to eighth order through triple-jump factorisation. A multilevel block-time-step refinement decreases the step size only for the innermost or most strongly-coupled stars, while maintaining symplecticity, and a binary-tree tiling lets non-overlapping pairs update simultaneously, giving $\mathcal{O}(N^{2}/P)+2P$ scaling on $P$ processors. The result is an exactly orbital-energy-conserving and angular-momentum-vector conserving integrator whose dominant cost is $N^{2}$ pair rotations, but whose accuracy can be tuned by the time-step hierarchy and the chosen operator-splitting order. Details on N-Ring may be found in Kocsis and Tremaine (2015).